# Do Larger Language Models Imply Better Generalization? A Pretraining Scaling Law for Implicit Reasoning

### v2

@@ 5.3 Limitations
We want to highlight that this study is only conducted on simplified pretraining data from knowledge graphs, and the results are likely not directly applicable to real-world language model pretraining with large text corpus. The setting of our study provides a reasonable analogy to the real-world language model pretraining, and the obtained insight might be found useful in the real world when the compute is abundant with very large models and very large datasets that exhaustively traverse the underlying knowledge graph. We leave the work of verifying our scaling law in the real word to future research due to its resource-demanding nature.
@@ 6 Related Work
Our paper focuses on the reasoning capability of language models which has drawn a lot of attention recently (Zhang et al., 2023; Chen et al., 2023; Yao et al., 2023a; b; Wang et al., 2023; Guo et al., 2025; Jin et al., 2024; Yeo et al., 2025; Team et al., 2025; Li et al., 2025). LLMs usually reason in a step-by-step manner in real-world tasks like math word problems (Wei et al., 2022b). In our experiments, we do not ask language models to generate a step-by-step solution for its answer, but ask the language model to directly choose the correct answer from the given options, because our pretrain-only language models are not trained to give a step-by-step solution for a query. Our synthetic reasoning environment is the most similar to Wang et al. (2024b), which also use the knowledge graph completion task as a testbed to understand how language models learn to reason at pretraining time. They propose that language models are able to aggregate random walk paths sampled from the knowledge graph. Wang et al. (2024a); Zhu et al. (2024) also employ a graph structure to ground their synthetic reasoning tasks to explain how LLMs reason, but their reasoning is defined as concatenations of relations: A is $r_{1}$ to B and B is $r_{2}$ to C implies A is $r_{1}r_{2}$ to C. The knowledge graph completion task we employ is more complex than simple concatenation of relations as the language model needs to find out which relation $r_{1}r_{2}$ corresponds to from the knowledge graph.
@@ 7 Conclusion
This paper investigates reasoning scaling in language models trained on knowledge graphs. Our results reveal a U-shaped relationship between model size and reasoning performance, where overparameterization leads to excessive memorization and degraded reasoning ability. We identify key factors that determine the optimal model size, such as the number of training triples and graph complexity. Notably, we propose an empirical scaling law linking optimal model size to graph search entropy, offering a quantitative guide for model design. While our experiments are conducted in controlled settings, these insights pave the way for future work in real-world pretraining scenarios and improved reasoning capabilities in LLMs.
Appendix A Experiment Details
@@
Table 3: Knowledge graph hyperparameter settings for Figure 3 experiments. We keep $L_{min}=2$ and $L_{max}=4$ for all experiments. Here $N$ denotes the number of triples, $N_{e}$ denotes the number of entities, $N_{r}$ denotes the number of relations, $N_{h}$ denotes the number of rules, $\gamma$ denotes the ratio between deductible triples and atomic triples, $L_{min}$ denotes the minimum rule length, and $L_{max}$ denotes the maximum rule length.
Appendix B Synthetic Knowledge Graph Generation Code

### v3

@@ 5.3 Limitations
We want to highlight that this study is only conducted on simplified pretraining data from knowledge graphs, and the results are not directly applicable to real-world language model pretraining with large text corpus. The setting of our study provides a reasonable analogy to the real-world language model pretraining, and the obtained insight might be found useful in the real world when the compute is abundant with very large models and very large datasets that exhaustively traverse the underlying knowledge graph. We leave the work of verifying our scaling law in the real word to future research due to its resource-demanding nature.
@@ 6 Related Work
Our paper focuses on the reasoning capability of LMs which has drawn a lot of attention recently (Zhang et al., 2023; Chen et al., 2023; Yao et al., 2023a; b; Wang et al., 2023; Guo et al., 2025; Jin et al., 2024; Yeo et al., 2025; Team et al., 2025; Li et al., 2025). LLMs are usually trained to reason in a step-by-step manner in real-world tasks like math problems (Wei et al., 2022b) and coding (Yang et al., 2024). In our experiments, we do not ask LMs to generate a CoT solution, but ask the language model to directly choose the correct answer from the given options, because our pretrain-only LMs are not trained to give a CoT solution for a query. Our synthetic reasoning environment is the most similar to Wang et al. (2024b), which also use the knowledge graph completion task as a testbed to understand how LMs learn to reason at pretraining time. They propose that LMs are able to aggregate random walk paths sampled from the knowledge graph. Wang et al. (2024a); Zhu et al. (2024) also employ a graph structure to ground their synthetic reasoning tasks to explain how LLMs reason, but their reasoning is defined as concatenations of relations: A is $r_{1}$ to B and B is $r_{2}$ to C implies A is $r_{1}r_{2}$ to C. The knowledge graph completion task we employ is more complex than simple concatenation of relations as the language model needs to find out which relation $r_{1}r_{2}$ corresponds to from the knowledge graph.
@@ 7 Conclusion
This paper presents a rigorous study of the scaling behavior of implicit reasoning in language models pretrained on knowledge graphs. Our findings reveal a U-shaped relationship between implicit reasoning performance and model size: overparameterization induces excessive memorization, which in turn degrades reasoning ability. We further identify key factors that determine the optimal model size, including the number of training triples and the complexity of the graph. Most notably, we propose an empirical scaling law that links the optimal model size to graph search entropy, demonstrating that a language model can reason over approximately 0.008 bits of information per parameter. Although our experiments are conducted in controlled settings to ensure rigor, the insights derived from this work offer promising directions for future studies on real-world pretraining and the enhancement of reasoning capabilities in large language models.
8 Reproducibility statement
We have taken several steps to ensure the reproducibility of our results. All experimental settings, including model architectures, training procedures, and hyperparameters, are described in detail in Section 2, Section 5 and Appendix B. To facilitate empirical reproducibility, we include a script of data construction steps in Appendix D. Additionally, we provide a simplified version of our source code as an easy-to-run Jupyter notebook for reproducing some of our experiments in the supplementary materials.
Appendix
Appendix A Data Processing Example
Figure 5: An example of a triple being processed in three different ways.
Appendix B Experiment Details
@@
Table 3: Knowledge graph hyperparameter settings for Figure 3 experiments. We keep $L_{min}=2$ and $L_{max}=4$ for all experiments. Here NN denotes the number of triples, $N_{e}$ denotes the number of entities, $N_{r}$ denotes the number of relations, $N_{h}$ denotes the number of rules, $\gamma$ denotes the ratio between deductible triples and atomic triples, $L_{min}$ denotes the minimum rule length, and $L_{max}$ denotes the maximum rule length.
Appendix C Synthetic Knowledge Graph v.s. Real-world Knowledge Graph
Figure 6: Distribution of node degrees of synthetic and real-world knowledge graphs.
Figure 7: Distribution of number of outgoing relations per node of synthetic and real-world knowledge graphs.
Appendix D Synthetic Knowledge Graph Generation Code